# Thompson sampling MAB multi-armed bandit

### v2

@@ VI Case study: call option trading
Figure 3: Flow Chart for call option and trade execution using SHVV and UCB-VV.
Top 8 Stocks $49$ $48$ $27$ $62$ $46$ $57$ $42$ $41$
Rewards $195.53$ $160.63$ $843.53$ $54.23$ $264.81$ $616.77$ $161.61$ $31.66$
Total Reward $2328.81$
Total Premium $622.74$
Table I: Top 8 stocks with the reward accumulated by option trading and the premium paid.
[…]
Experiment: We employ GBM to simulate stock prices, a widely adopted stochastic process in quantitative finance for its ability to model realistic stock price trajectories, making it invaluable for back-testing trading strategies, evaluating investment opportunities, and pricing financial derivatives. To ensure the analysis remains responsive to evolving market dynamics, a rolling window approach should be incorporated. This method prioritizes recent data, enabling the capture of emerging trends and volatility patterns that would otherwise be obscured or averaged out over longer time frames. By continuously updating the dataset to reflect the most current market conditions, the rolling window enhances both the relevance and accuracy of the analysis. For this study, we also used a rolling window of size $\tau=90$. This dynamic adaptation makes it especially valuable for time-sensitive financial applications, where timely insights are critical for informed decision-making.
@@
We implement SHVV algorithm to shortlist “top 8” stocks referring to the $8$ stocks with the highest variances given a budget of $n_{1}=20e3$, thus pulling each arm $1674$ times. We then employ UCB-VV to select a stock for trading at time $t$ with an additional budget of $n_{2}=1000$. If the condition $S_{t}>R$ is met, we collect the corresponding profits. This process continues iteratively: at each time step, we select the next stock according to the UCB-VV index, execute the trade if the condition holds, and accumulate the profits across all selected stocks until time horizon $n=n_{1}+n_{2}$ as shown in Table I.
We further compare the performance of our proposed framework with UCB for $n=1000$. We see in Fig 4, for $n=1000$, we achieve substantially higher cumulative rewards compared to UCB.
@@
Figure 4: Cumulative reward (profit) v/s trading period of GBM simulated $100$ stocks for the proposed strategy and UCB.
VII Appendices
VII-A Proof of Theorem 1
@@
We need to bound each $s_{i}(n)$ to bound the regret. Let $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$ be upper confidence bound. Let $z$ be any arbitrary positive integer and assume that the arm $i$ has been pulled $z$ times, including the pilot fraction. We use a subscript * to any quantity that refers to the optimal arm. At the start of the experiment, each arm is pulled at least once since $S_{\rm P}=K/n$, including the sub-optimal arms. Thus,
@@
$\sigma^{2}_{*}<\sigma_{i}^{2}+2C_{t,s_{i}}$ (8)
$C_{t,s_{*}}$ and $C_{t,s_{i}}$ are the bias terms for the optimal and sub-optimal arm, respectively. We bound the probability of events (6) and (7) using Lemma 1:
@@
Similarly for $C_{t,s_{i}}$, probability of (7) will be $\leq t^{-4}$. We see that the condition in equation (8) arises when the true variances of the optimal arm and sub-optimal arm are close to each other. If the true variance of optimal is less than the true variance of any sub-optimal arm with the addition of 2 times the uncertainty term, i.e., $C_{t,s_{i}(q)}$, the agent is bound to make an error until enough number of pulls have taken place. As $z$ represents the initial number of pulls, and if pulls are sufficient, then equation (8) becomes false. We see that for $z=\Big{\lceil}\frac{8\log{t}}{\delta^{2}_{i}}\Big{\rceil}$, i.e., $s_{i}=z$, the number of pulls of the sub-optimal arm is substantial enough to make the occurrence of an error event to become highly unlikely. This can be seen as,
@@
$\geq\sigma^{2}_{*}-\sigma_{i}^{2}-2\sqrt{\frac{2\log{t}\delta^{2}_{i}}{8\log{t}}}=\sigma^{2}_{*}-\sigma_{i}^{2}-\delta_{i}=0.$
@@
$\text{R}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
Note that for $s_{i}\geq{\frac{8\log{n}}{\delta^{2}_{i}}}$, the probability of error diminishes at a rate that contributes to logarithmic regret.
VII-B Proof of Lemma 2
@@
[…]
@@
[…]
@@ VII-C Proof of Theorem 2
[…]
@@
[…]
$\implies\;\liminf_{n\rightarrow\infty}\frac{R(n)}{\log n}$ $\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}$
VII-D Proof of Lemma 4
@@ VII-E Proof of Theorem 4
A bounded distribution is defined on the interval $[l,u]$. The ${\rm KL}$ divergence between 2 such distributions $\nu$ and $\tilde{\nu}$ depends on their PDFs as given below:
[…]
Let $X_{i}(s)$ denote the samples drawn from arm $i$, which corresponds to arm in the first bandit setting $G^{1}$ The empirical ${\rm KL}$-divergence is given as:
@@ VII-F Proof of Theorem 5
$2\exp\left(-\frac{n\epsilon}{4v^{2}}\right).$
Thus, $\exists C:0<C<\min\left\{\frac{1}{4},C_{5}\right\}$ such that the statement of the theorem follows. ∎
VII-G Proof of Theorem 6
@@
$|\bar{S}(n)-S|=\left|\frac{\bar{\mu}(n)}{\bar{V}(n)}-\frac{\mu}{\sigma^{2}}\right|=\left|\frac{\bar{\mu}(n)\sigma^{2}-\mu\bar{V}(n)}{\bar{V}(n)\sigma^{2}}\right|.$
@@
$|\bar{S}(n)-S|\leq\frac{|\bar{\mu}(n)-\mu|}{\bar{V}(n)}+\frac{|\mu|\cdot|\bar{V}(n)-\sigma^{2}|}{\bar{V}(n)\sigma^{2}}.$
We first leverage concentration inequalities to stabilize the denominator by enforcing its constancy, i.e., to show with high probability that $\bar{V}(n)\geq\frac{\sigma^{2}}{2}$. Utilizing the result of Theorem 5, where we put $\epsilon=\frac{\sigma^{2}}{2}$ and get:
$\mathbb{P}\left(\bar{V}(n)\geq\frac{\sigma^{2}}{2}\right)\leq 2\exp\left(-Cn\min\left(\frac{\sigma^{4}}{4v^{4}},\frac{\sigma^{2}}{2v^{2}}\right)\right).$
Thus with probability at least $1-2\exp\left(-Cn\min\left(\frac{\sigma^{4}}{4v^{4}},\frac{\sigma^{2}}{2v^{2}}\right)\right)$, we have $\bar{V}(n)\geq\frac{\sigma^{2}}{2}\,\Rightarrow\,\frac{1}{\bar{V}(n)}\leq\frac{2}{\sigma^{2}}$. For this event, the deviation bound becomes:
$|\bar{S}(n)-S|\leq\frac{2|\bar{\mu}(n)-\mu|}{\sigma^{2}}+\frac{2|\mu|\cdot|\bar{V}(n)-\sigma^{2}|}{\sigma^{4}}.$
@@
$|\bar{\mu}(n)-\mu|\leq\frac{\eta\sigma^{2}}{2},\quad|\bar{V}(n)-\sigma^{2}|\leq\frac{\eta\sigma^{4}}{2|\mu|}.$
Using the concentration bound for the sample mean $\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\epsilon\right)\leq 2\exp\left(-\frac{n\epsilon^{2}}{2v^{2}}\right)$, we set $\epsilon=\frac{\eta\sigma^{2}}{2}$ to obtain:
$\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\frac{\eta\sigma^{2}}{2}\right)\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{4}}{8v^{2}}\right).$
Similarly, invoking the sub-exponential tail bound stated in Theorem 5 for the deviation $|\bar{V}(n)-\sigma^{2}|$ and setting $\epsilon_{1}=\frac{\eta\sigma^{4}}{4|\mu|}$, we obtain
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|\geq\frac{\eta\sigma^{4}}{2|\mu|}\right)\leq$
$\hskip 56.9055pt2\exp\left(-C_{6}n\min\left(\frac{\eta^{2}\sigma^{8}}{4\mu^{2}v^{4}},\frac{\eta\sigma^{4}}{2|\mu|v^{2}}\right)\right),$
for some constant $C_{6}>0$. Combining these bounds via the union bound, the total probability of deviation satisfies:
$\mathbb{P}\left(|\bar{S}(n)-S|\geq\eta\right)$ $\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{4}}{8v^{2}}\right)+$
$2\exp\left(-C_{6}n\min\left(\frac{\eta^{2}\sigma^{8}}{4\mu^{2}v^{4}},\frac{\eta\sigma^{4}}{2|\mu|v^{2}}\right)\right).$

### v3

@@ VI-A Relation Between Variance and Call Option Payoff
Figure 3: Flow Chart for call option and trade execution using SHVV and UCB-$V$.
Our setting is based on a straightforward premise: investors/buyers of call options favor “good uncertainty” as it enhances the potential for significant gains, but are averse to “bad uncertainty”, which raises the risk of substantial losses. We define “good uncertainty” as volatility linked to positive stock market returns and “bad uncertainty” as volatility tied to negative returns. The arm is played only if $S_{t}>R$ and the buyer profits. The formulation of a call option provides protection against negative returns. We neglect all transaction costs and taxes. Fig. 3 represents the flowchart for using UCB-VV and SHVV for call option trading of GBM simulated daily returns for $100$ stocks.
@@
Experiment: We employ GBM to simulate stock prices, a widely adopted stochastic process in quantitative finance for its ability to model realistic stock price trajectories, making it invaluable for back-testing trading strategies, evaluating investment opportunities, and pricing financial derivatives. We implement SHVV algorithm to shortlist “top 8” stocks referring to the $8$ stocks with the highest variances given a budget of $n_{1}=20e3$. We then employ UCB-VV to select a stock for trading at time $t$ with an additional budget of $n_{2}=1000$. If the condition $S_{t}>R$ is met, we collect the corresponding profits. This process continues iteratively: at each time step, we select the next stock according to the UCB-VV index, execute the trade if the condition holds, and accumulate the profits across all selected stocks until time horizon $n=n_{1}+n_{2}$. We compare the performance of our proposed framework with a two stage baseline i.e, SH-Mean followed by UCB1 under same budget restrictions. SH-Mean selects the top-8 arms with the highest empirical mean (analogous to SHVV). We see in Fig 2, SHVV + UCB-VV achieve substantially higher cumulative rewards compared to SH-Mean + UCB1.
@@
Figure 4: Cumulative profit of GBM simulated 100100 stocks for the proposed two stage strategy i.e, SHVV + UCB-VV v/s SH-Mean+ UCB1.
VII Future work
While this work focuses on the stochastic setting with stationary rewards, several promising extensions arise naturally from both theoretical and practical perspectives. First, a minimax formulation of variance bandits where the goal is to minimize misallocation uniformly over a class of distributions would be particularly relevant in adversarial or distribution-free environments where the variance gap is unknown or potentially vanishing. Second, real-world financial time series are inherently non-stationary, with volatility evolving dynamically over time. Extending UCB-VV and SHVV to non-stationary settings via sliding windows, change-point detection, or weighted estimators would significantly enhance their applicability to live trading systems. These directions not only align with the broader vision of robust, risk-aware decision-making but also open new avenues for bridging bandit theory with practical financial engineering.
VIII Appendices
VIII-A Proof of Theorem 1
@@
We need to bound each $s_{i}(n)$ to bound the misallocation. Let $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$ be upper confidence bound. Let $z$ be any arbitrary positive integer and assume that the arm $i$ has been pulled $z$ times, including the pilot fraction. We use a subscript * to any quantity that refers to the optimal arm. At the start of the experiment, each arm is pulled at least once since $S_{\rm P}=K/n$, including the sub-optimal arms. Thus,
@@
$\delta_{i}=\sigma_{*}^{2}-\sigma_{i}^{2}\leq C_{t-1,s_{i}}+C_{t-1,s_{*}},$ (8)
$C_{t,s_{*}}$ and $C_{t,s_{i}}$ are the bias terms for the optimal and sub-optimal arm, respectively.
We bound the probability of events (6) and (7) using Lemma 1:
@@
Similarly for $C_{t,s_{i}}$, probability of (7) will be $\leq t^{-4}$. We see that the condition in equation (8) arises when the true variances of the optimal arm and sub-optimal arm are close to each other. Then, the agent is bound to make an error until enough number of pulls have taken place. We split the analysis into two cases:
  • Case 1: $s_{*}\geq z$; in this case, $C_{t,s_{*}}\leq C_{t,z}$, and we bound $C_{t,s_{i}}+C_{t,s_{*}}\leq 2C_{t,z}$.
  • Case 2: $s_{*}<z$; this implies the optimal arm has been pulled fewer than $z$ times. But since the algorithm includes an initial exploration phase (each arm pulled at least once) and $s_{*}$ grows over time, this case can only occur for a limited number of time steps. We bound the total contribution of this case by a constant.
By choosing $z=\left\lceil\frac{8\log n}{\delta_{i}^{2}}\right\rceil$, we ensure that $C_{t,s_{i}}+C_{t,s_{*}}<\delta_{i}$. Thus, for $z=\Big\lceil\frac{8\log{t}}{\delta^{2}_{i}}\Big\rceil$, i.e., $s_{i}=z$, the number of pulls of the sub-optimal arm is substantial enough to make the occurrence of an error event to become highly unlikely. This can be seen as,
@@
$\hskip-56.9055pt\geq\sigma^{2}_{*}-\sigma_{i}^{2}-2\sqrt{\frac{2\log{t}\delta^{2}_{i}}{8\log{t}}}=\sigma^{2}_{*}-\sigma_{i}^{2}-\delta_{i}=0.$
@@
$\mathcal{M}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}^{2}}+1+\frac{\pi^{2}}{3}.$
Note that for $s_{i}\geq{\frac{8\log{n}}{\delta^{2}_{i}}}$, the probability of error diminishes at a rate that contributes to logarithmic misallocation.
VIII-B Proof of Lemma 2
@@
$\text{Also,}\;\;\mathbb{P}_{\mathcal{F}}[S(n)]=\mathbb{E}_{\mathcal{F}^{i}}[\mathbb{I}_{S(n)}e^{\gamma}],$ (14)
@@
$\text{Equivalently,}\;\mathbb{P}_{\mathcal{F}}\bigg[s_{i}\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg]\rightarrow 1\,\text{as}\,n\rightarrow\infty$
@@ VIII-C Proof of Theorem 2
$\text{So,}\;\;\lim_{n\rightarrow\infty}\frac{\mathbb{E}[s_{i}]}{\log n}$ $\geq\lim_{n\rightarrow\infty}\mathbb{P}\left[s_{i}\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\right]\frac{C_{1}}{I(f_{i},f_{*})}$
@@
$\text{We know,}\;\;\mathcal{M}(n)=\sum_{i}^{K}\mathbb{E}[s_{i}]$ $\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\log n}{I(f_{i},f_{*})}$
$\implies\;\liminf_{n\rightarrow\infty}\frac{\mathcal{M}(n)}{\log n}$ $\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}}{I(f_{i},f_{*})}$
Therefore, the lower bound is order-optimal.
VIII-D Proof of Lemma 4
@@ VIII-E Proof of Theorem 4
A bounded distribution is defined on the interval $[l,u]$. The $\rm KL$} divergence between 2 such distributions $\nu$ and $\tilde{\nu}$ depends on their PDFs as $I(\nu,\tilde{\nu})=\int_{l}^{u}\log\left(\frac{f_{\nu}(x)}{f_{\tilde{\nu}}(x)}\right)f_{\nu}(x){\rm d}x$. Let $X_{i}(s)$ denote the samples drawn from arm $i$, which corresponds to arm in the first bandit setting $G^{1}$. The empirical $\rm KL$}-divergence is given as:
@@ VIII-F Proof of Theorem 5
$\hskip 85.35826pt2\exp\left(-\frac{n\epsilon}{4v^{2}}\right)$ (21)
Thus, $\exists C_{V}:0<C_{V}<\min\left\{\frac{1}{4},C_{5}\right\}$ such that the statement of the theorem follows. ∎
VIII-G Proof of Theorem 6
@@
$|\bar{S}(n)-S|=\left|\frac{\bar{\mu}(n)}{\sqrt{\bar{V}(n)}}-\frac{\mu}{\sigma}\right|=\left|\frac{\bar{\mu}(n)\sigma-\mu\sqrt{\bar{V}(n)}}{\sqrt{\bar{V}(n)}\sigma}\right|.$
@@
$|\bar{S}(n)-S|\leq\frac{|\bar{\mu}(n)-\mu|}{\sqrt{\bar{V}(n)}}+\frac{|\mu|\cdot|\sqrt{\bar{V}(n)}-\sigma|}{\sqrt{\bar{V}(n)}\sigma}.$
We first leverage concentration inequalities to stabilize the denominator by enforcing its constancy, i.e., to show with high probability that $\bar{V}(n)\geq\frac{\sigma^{2}}{4}$. Utilizing the result of Corollary 1 we get:
$\mathbb{P}\left(\sqrt{\bar{V}(n)}\geq\frac{\sigma}{2}\right)\leq 2\exp\left(-Cn\min\left(\frac{\sigma^{2}}{2v^{4}},\frac{\sigma}{2v^{2}}\right)\right).$
Thus with probability at least $1-2\exp\left(-Cn\min\left(\frac{\sigma^{2}}{2v^{4}},\frac{\sigma}{2v^{2}}\right)\right)$, we have $\sqrt{\bar{V}(n)}\geq\frac{\sigma}{2}\,\Rightarrow\,\frac{1}{\sqrt{\bar{V}(n)}}\leq\frac{2}{\sigma}$. For this event, the deviation bound becomes:
$|\bar{S}(n)-S|\leq\frac{2|\bar{\mu}(n)-\mu|}{\sigma}+\frac{2|\mu|\cdot|\sqrt{\bar{V}(n)}-\sigma|}{\sigma^{2}}.$
@@
$|\bar{\mu}(n)-\mu|\leq\frac{\eta\sigma}{4},\quad|\sqrt{\bar{V}(n)}-\sigma|\leq\frac{\eta\sigma^{2}}{4|\mu|}.$
Using the concentration bound for the sample mean $\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\epsilon\right)\leq 2\exp\left(-\frac{n\epsilon^{2}}{2v^{2}}\right)$, we set $\epsilon=\frac{\eta\sigma}{4}$ to obtain:
$\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\frac{\eta\sigma}{4}\right)\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{2}}{32v^{2}}\right).$
Similarly, invoking the sub-exponential tail bound stated in Corollary 1 and setting $\epsilon_{1}=\frac{\eta\sigma^{2}}{4|\mu|}$, we obtain
$\mathbb{P}\left(|\sqrt{\bar{V}(n)}-\sigma|\geq\frac{\eta\sigma^{2}}{4|\mu|}\right)\leq$
$\hskip 56.9055pt2\exp\left(-C_{7}n\min\left(\frac{\eta^{2}\sigma^{4}}{16|\mu|^{2}v^{4}},\frac{\eta\sigma^{2}}{4|\mu|v^{2}}\right)\right),$
for some constant $C_{7}>0$. Combining these bounds via the union bound, the total probability of deviation satisfies:
$\mathbb{P}\left(|\bar{S}(n)-S|\geq\eta\right)$ $\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{2}}{32v^{2}}\right)+$
$2\exp\left(-C_{7}n\min\left(\frac{\eta^{2}\sigma^{4}}{16|\mu|^{2}v^{4}},\frac{\eta\sigma^{2}}{4|\mu|v^{2}}\right)\right).$